OpenStudy (anonymous):

find second derivative of e^(8x) +3cos(x)sin(x)

OpenStudy (anonymous):

\[e ^{8x} +3\sin(x)\cos(x)\]

OpenStudy (anonymous):

for the first derivative I got \[8e ^{8x} +3\cos^2x-3\sin^2x\] is that right so far?

OpenStudy (anonymous):

Yes, I believe so, your first derivative looks correct.

OpenStudy (anonymous):

but then I get confused about how to get the derivative of that

OpenStudy (anonymous):

Well, we take it in steps, the first part is just \[64e ^{8x}\]Second and third portion use the chain rule. Second portion:\[-6cosx*sinx\] and third portion: \[-6sinx*cosx\]

OpenStudy (anonymous):

Do you need help understanding the chain rule?